OpenStudy (anonymous):

what is the slope and the Y-intercept of this line: Y=-2x+8

OpenStudy (anonymous):

\[\large \sf y=mx+b\]\[\large \sf m=slope\]\[\large \sf b=Y-intercept\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Sorry, didn't see the 'x' in front of -2 You could also set 'x' to 0 and solve for 'Y' to get the intercept. -2*0=0 y=0+8 0+8=8 y=8

OpenStudy (anonymous):

Thank you for the help

OpenStudy (anonymous):

Np.